# DP-LDMs: Differentially Private Latent Diffusion Models

### v5

@@ A.2.1 Choosing public dataset with DP constraint
. $\epsilon$ 0.1 0.5 1 2 5 10 SVHN 78.9930 17.3465 8.9223 4.7426 2.1582 1.3368 KMNIST 79.9224 17.5794 9.0605 4.6092 1.9300 1.0794 EMNIST 80.4267 17.3617 8.9058 4.4561 1.8926 1.0242 $\epsilon$ 0.1 0.5 1 2 SVHN 0.2675 0.2662 0.2663 0.2663 KMNIST 0.0489 0.0459 0.0457 0.0457 EMNIST 0.0228 0.0201 0.0201 0.0197
@@ A.4 Transferring from Imagenet to CelebA32
We also apply our model in the task of generating $32\times 32$ CelebA images. The same pretrained autoencoder as our CIFAR-10 experiments in Section 5.1 is used, but since this experiment is for unconditional generation, we are unable to re-use the LDM. A new LDM is pretrained on Imagenet without class conditioning information, and then fine-tuned on CelebA images scaled and cropped to $32\times 32$ resolution. Our FID results for $\delta=10^{-6}$, $\epsilon=1,5,10$ are summarized in Table 16. We achieve similar results to DP-MEPF for $\epsilon=5$ and $\epsilon=10$. As with our results at $64\times 64$ resolution, our LDM model does not perform as well in higher privacy settings ($\epsilon=1$). Sample images are provided in Figure 10

### v6

@@ 6 Conclusion
Broader Impact Statement
As investigated in Carlini et al. (2023), diffusion models can memorize individual images from the training data set and generate an identical synthetic image to the training image. Aiming to impact society positively, we provide a method to fine-tune latent diffusion models with differential privacy guarantees so that the fine-tuned models do not generate identical images to those in the private data.
Our method relies on public data for better scalability of differential privacy, which needs some attention. As Tramèr et al. (2022) pointed out, public data themselves may still contain sensitive information. From our perspective and as many other DP generative modelling papers also noticed, auxiliary public data still emerges as the most promising option for attaining satisfactory utility, for the models at this large scale. We hope for better-curated public datasets to be available in the near future.
During the review process, one of the reviewers mentioned that generative tools are becoming more and more realistic and of general availability, with the risk of generating harmful content, especially when generating human-related content. Broadly speaking, our method does not stop malicious users from creating harmful content, as the diffusion model will generate images based on any input prompt given by the malicious user. However, what our method can do is protect the privacy of the individuals whose data was in the private dataset, meaning the faces in the private dataset will not appear when generating images from the DP fine-tuned models.
To encourage reproducible research practices, our code is available at https://github.com/ParkLabML/DP-LDM with detailed instructions. And all the hyperparameters are discussed in detail in Appendix Sec. B.
Limitations
While the paper proposes a technique to counter-attack the privacy risks of diffusion models, the evaluation has not been done against existing privacy attacks. So we know that the diffusion model is private “under the definition of DP", but we caution practitioners against making an assumption that a DP image generative model protects private data against all possible attack methods.
This may not be a direct limitation of our work itself, but it is restrictive that there is no agreement within the community on which definition of privacy should be used to assess the privacy of generated images. Carlini et al. (2023) provided a single example of exploitation, but it remains unclear what aspects we aim to protect in generated data. Unfortunately, the privacy parameters $\epsilon$ and $\delta$ in our method are especially difficult to interpret in the image domain. While private images may be protected, it is unclear whether perceptually similar images enjoy the same level of privacy. Further work is required to suggest more appropriate definitions of privacy for practitioners to employ to avoid providing a vague sense of privacy in the image generation domain.
Acknowledgments
We thank our anonymous reviewers for their constructive feedback, which has helped significantly improve our manuscript. We thank the Digital Research Alliance of Canada (Compute Canada) for its computational resources and services. M. Liu was funded by the Canada Graduate Scholarships — Master’s program of the Natural Sciences and Engineering Research Council of Canada (NSERC). S. Lyu, M. Vinaroz, and M. Park were supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Canada CIFAR AI Chairs program. M. Park was also partially funded by Novo Nordisk Fonden RECUIT grant no.0065800 during her stay at the Technical University of Denmark.
@@ A.2.1 Choosing public dataset with DP constraint
$\epsilon$ 0.1 0.5 1 2 5 10
SVHN 78.9930 17.3465 8.9223 4.7426 2.1582 1.3368
KMNIST 79.9224 17.5794 9.0605 4.6092 1.9300 1.0794
EMNIST 80.4267 17.3617 8.9058 4.4561 1.8926 1.0242
$\epsilon$ 0.1 0.5 1 2
SVHN 0.2675 0.2662 0.2663 0.2663
KMNIST 0.0489 0.0459 0.0457 0.0457
EMNIST 0.0228 0.0201 0.0201 0.0197
@@ A.4 Transferring from Imagenet to CelebA32
We also apply our model in the task of generating $32\times 32$ CelebA images. The same pretrained autoencoder as our CIFAR-10 experiments in Section 5.1 is used, but since this experiment is for unconditional generation, we are unable to re-use the LDM. A new LDM is pretrained on Imagenet without class conditioning information, and then fine-tuned on CelebA images scaled and cropped to $32\times 32$ resolution. Our FID results for $\delta=10^{-6}$, $\epsilon=1,5,10$ are summarized in Table 16. We achieve similar results to DP-MEPF for $\epsilon=5$ and $\epsilon=10$. As with our results at $64\times 64$ resolution, our LDM model does not perform as well in higher privacy settings ($\epsilon=1$). Sample images are provided in Fig. 10
@@ Appendix C Additional Samples
Appendix D Comparison between DP and non-DP Samples
Figure 13: Synthetic CelebAHQ images from a non-DP-trained LDM (Top, black box, images taken from Rombach et al. (2022)) versus those from DP-trained LDMs (Bottom, red box). The non-DP synthetic images exhibit humanly discernible characteristics of faces. On the other hand, DP synthetic samples exhibit distortions at varying levels, where a higher privacy guarantee ($\epsilon=1$) yields more distortion.